OpenStudy (anonymous):

I'm having trouble with a definite integral that involves trigonometric substitution. I was wondering if I could receive some advice on how to work it.

OpenStudy (anonymous):

\[\int\limits_{0}^{4} (64-x^{2})^{-2}\]

OpenStudy (anonymous):

So for this integral, you have a term which looks like: a^2-x^2 This is the form where you make a sine substitution. Just for note, the other possible forms: - You would use a tangent substitution for a^2+x^2 - You would use a secant substitution for x^2 - a^2 So in this case, we use sine. Draw a triangle: So here, we can see in the drawing:\[\sin(\theta)=\frac{x}{a}\]This gives us our 'x' and 'dx':\[x = asin(\theta)\]\[dx = acos(\theta)d \theta\] We now adapt this to the integral.

OpenStudy (anonymous):

Sorry having connection issues

OpenStudy (anonymous):

Okay, so what we need to do is re-draw our triangle with our specific values: